OpenStudy (anonymous):

Write the equation in slope-intercept form. What are the slope and y-intercept? -11x + 9y = -12 A. y = 11/9x - 4/3 slope: 11/9 y-intercept: -4/3 B. y = 11/9x - 4/3 slope: 11/9 y-intercept: 4/3 C. y = -11/9x + 4/3 slope: 11/9 y-intercept: 4/3 D. y = 11/9x + 4/3 slope: 4/3 y-intercept: 11/9

OpenStudy (anonymous):

Have you tried solving for 'y'?

OpenStudy (anonymous):

the equation of a line (generally) is as follows: y=mx+b where m is the slope and b is the y-intercept.

OpenStudy (anonymous):

can you explain better i really dont understand

OpenStudy (anonymous):

Okay, so basically you want to get whatever equation you have into something that has y= some function of x. do you agree?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay good, so now you solve for y in the equation that was given to you; which was -11x + 9y = -12

OpenStudy (anonymous):

so you want to get y by itself

OpenStudy (anonymous):

dont know what i did but i got -1.2x+1.3=y.........

OpenStudy (anonymous):

your numbers are right (but try leaving it in fraction form), however your signs are off

OpenStudy (anonymous):

would 1.3 be negative and 1.2 be positive

OpenStudy (anonymous):

correct. so your answer should be: y=11/9x - 4/3

OpenStudy (anonymous):

thank you:)

OpenStudy (anonymous):

no problem! :) and thanks for the medal!

OpenStudy (anonymous):

no problem